OpenStudy (anonymous):

A scientist is studying the growth of a particular species of plant. He writes the following equation to show the height of the plant f(n), in cm, after n days: f(n) = 15(1.04)n Part A: When the scientist concluded his study, the height of the plant was approximately 20.53 cm. What is a reasonable domain to plot the growth function? Part B: What does the y-intercept of the graph of the function f(n) represent? Part C: What is the average rate of change of the function f(n) from n = 1 to n = 5, and what does it represent?

OpenStudy (anonymous):

I know what Part B is. i just need help with part A+B

OpenStudy (anonymous):

I meant C

Directrix (directrix):

Is this >> f(n) = 15(1.04)n supposed to be f(n) = 15(1.04) ^n

OpenStudy (anonymous):

Yes

Directrix (directrix):

I can't help with part C until I know the results of parts A and B.

OpenStudy (anonymous):

For part B I put down that it represents the plant height

Directrix (directrix):

Read this and start at the section where @dan815 begins his work: http://openstudy.com/updates/53aaf4d3e4b02c45b013c23d

OpenStudy (anonymous):

Okay

Directrix (directrix):

Thatis a different problem but notice that the function is exponential. is f(n)= 8(1.05)^n . So, I'm asking you to check your equation.

OpenStudy (anonymous):

Alright

OpenStudy (anonymous):

Thanks